OpenStudy (anonymous):

Find the product of (4x + 3y)2. 16x2 - 24x + 9y2 16x2 + 24xy + 9y2 16x2 - 9y2 16x2 + 9y2

OpenStudy (texaschic101):

(4x + 3y)^2 = (4x + 3y)(4x + 3y) = 16x^2 + 12xy + 12xy + 9y^2 = 16x^2 + 24xy + 9y^2

OpenStudy (texaschic101):

do you have any questions ? I used the FOIL method to solve this problem

OpenStudy (anonymous):

I think I understand it

OpenStudy (anonymous):

thanks :)

OpenStudy (texaschic101):

anytime :)